Budgeted overhead for Sunland Company at normal capacity of 30000 direct labor hours is $6 per hour variable and $4 per hour fixed. In May, $302400 of overhead was incurred in working 31500 hours when 32000 standard hours were allowed.

The overhead controllable variance is

a) $5000 favorable

b) $17600 favorable

c) $9600 favorable

d) $17600 unfavorable
Answer #1

Calculation of overhead controllable variance:

Actual overhead= $302400

Budgeted fixed expense= 30000*4= $120000

Budgeted variable expense= 32000*6= 192000

Overhead controllable variance= Actual overhead-(budgeted fixed expense+variable budgeted expense)

Overhead controllable variance= 302400-(120000+192000)

Overhead controllable variance= -9600

So correct answer is c)$9600 favorable